Bergan Company estimates that total factory overhead costs will be $1,539,000 for the year. Direct labor hours are estimated to be 513,000. Required: A. For Bergan Company, determine the predetermined factory overhead rate using direct labor hours as the activity base. Round your answer to the nearest cent. B. During May, Bergan Company accumulated 19,000 hours of direct labor costs on Job 200 and 23,000 hours on Job 305. Determine the amount of factory overhead applied to Jobs 200 and 305 in May. C. Prepare the journal entry on May 30 to apply factory overhead to both jobs in May according to the predetermined overhead rate. Refer to the Chart of Accounts for exact wording of account titles.

A. For Bergan Company, determine the predetermined factory overhead rate using direct labor hours as the activity base. Round your answer to the nearest cent.

per direct labor hour

B. During May, Bergan Company accumulated 19,000 hours of direct labor costs on Job 200 and 23,000 hours on Job 305. Determine the amount of factory overhead applied to Jobs 200 and 305 in May.
Answer #1

a) Predetermine factory overhead rate = estimated overhead/estimated labor hour = 1539000/513000 = 3 per labor hour

b) Overhead applied:

Job 200 = 19000*3 = 57000

Job 305 = 23000*3 = 69000

c) Journal entry

Date account and explanation Debit Credit
Work in process-Job 200 57000
Factory overhead 57000
Date account and explanation Debit Credit
Work in process-Job 305 69000
Factory overhead 69000